OpenStudy (anonymous):

find the equationmof the line. write answer min slope - intercept form.. (-9,3) perendicular to -8x + 7y=51 Please and thank you! :)

OpenStudy (anonymous):

Thank you..

OpenStudy (anonymous):

7y=8x+51 y=-8/7x-7.29

OpenStudy (anonymous):

well heck.. i have 3 different answers. :(

OpenStudy (anonymous):

equation (x,y)(-9,3)=-8/7 change in y / change in x y-3/x+9=-8/7 7(y-3)=-8(x+9) 7y-21=-8x-72 7y=-8x-51 y=-8/7x-51/7 sorry for the mistake earlier the gradient -8/7 since the gradient for the first equation is 8/7 8/7*-8/7=-1

OpenStudy (anonymous):

did u get it now

OpenStudy (anonymous):

did u get it?